OpenStudy (love_to_love_you):

Please help, will fan and medal!!! Refer to the Concept Byte linear programming activity. Follow the step-by-step instructions at the top of the page. Complete each of the following steps for exercises # 1 & 3. 1. Rewrite the constraints in slope-intercept form. 2. List all vertices of the feasible region as ordered pairs. 3. List the values of the objective function for each vertex. 4. List the maximum or minimum amount, including the x, and y-value, of the objective function.

OpenStudy (gorv):

sorry online calculator is not working n by normal calc it is not possibble :(

OpenStudy (love_to_love_you):

I have to solve all odd problems.

OpenStudy (gorv):

which on e??

OpenStudy (love_to_love_you):

3, 5, 7, and 9

OpenStudy (love_to_love_you):

whatever you can help me with will be appreciated.

OpenStudy (gorv):

equation has no solution mean inconsistent

OpenStudy (love_to_love_you):

?

OpenStudy (gorv):

there are infinitely many solutions and the system is called dependent.

OpenStudy (gorv):

cost of 8 pen=8*0.69=5.52

OpenStudy (gorv):

let we purcahse x highlighter

OpenStudy (gorv):

cost=x*1.09 x*1.09+5.52=7.52 subtract 5.52from both side

OpenStudy (gorv):

x*1.09=2 divide both side by 1.09 x=2/1.09=1.8 which means we just purchased 1 highlighter

OpenStudy (love_to_love_you):

for number 3 it is an independent system right?

OpenStudy (gorv):

3 independent 4 dependent 5dependent 6 dependent 8 independet

OpenStudy (love_to_love_you):

I have to show work, did you just find that by looking at it or did you do anything to find that out?

OpenStudy (love_to_love_you):

5 is inconsistent

OpenStudy (love_to_love_you):

@gorv

OpenStudy (gorv):

5 has infinite solution

OpenStudy (love_to_love_you):

I literally have the answer sheet to check my work but I have to submit my work to my teacher.